OpenStudy (xxdragoonxx):

MEDAL! 1) If a train travels at an average speed of 37 miles per hour, how much time will it take to go 814 miles? Show your work.

OpenStudy (matlee):

814/37

OpenStudy (jhannybean):

\[d = r\cdot t\] where r = rate, t = time, and d = distance. We're given that r = 37 mi/hr , d = 814 miles. rearrange your function to solve for time. \[ \frac{rt}{r} = \frac{d}{r}\]\[ t = \frac{d}{r}\] Now just plug in your values. \[t = \frac{814 \ mi.}{37 \frac{mi}{hr}}= \frac{814}{37} \ mi\cdot \frac{hr}{mi} = \frac{814}{37} hr\]

OpenStudy (matlee):

-cough- 814/37

OpenStudy (jhannybean):

Units, my dear friend.

OpenStudy (matlee):

i see